# Revisiting boundary-driven method for transport: Finite-size effects and the role of system-bath coupling

### v1

@@ V.1 Integrable system
For this subsection, we consider the spin-1/2 XXZ chain with open boundary conditions described by Eq. (1).
In Fig. 2 the real-time dynamics of (a) the local magnetization density $\langle S_{r}^{z}\rangle$ and (b) the local current $\langle j_{r}\rangle$ for a spin-$1/2$ XXZ chain with $L=20$ sites and anisotropy $\Delta=1.5$ under a strong system-bath coupling $\gamma/J=1.0$ and weak driving $\mu=0.1$ is shown. Data obtained by stochastic unraveling (SU) is averaged over $N_{\text{traj}}=50000$ trajectories. The results of both methods, SU and TEBD, show a reasonable agreement. While the remaining deviation stems from the approximative nature of the two methods (finite number of trajectories/finite bond dimension), this deviation does not affect the conclusions drawn later. As can be seen from Fig. 2, panel (a) and (b), at times $t\gtrsim 100$, the resulting NESS is reached. The NESS features a linear magnetization profile in the bulk of the system and a NESS current independent of site location $r$. As a concrete example, in Fig. 2(c), we show the corresponding magnetization profile for a chosen time $tJ=200$. By fitting over the bulk sites $r=5\dots 15$, we can extract a linear slope of $\nabla S_{\text{ness}}^{z}=-0.0421\mu/2$, which is also consistent with the TEBD calculations. In Fig. 2(b) the local current according to Eq. (7) for the bulk sites $r=5\dots 15$ is shown. As for the magnetization profile, also here the choice of the initial infinite-temperature state is reflected in $\langle j_{r}(t=0)\rangle=0$. For times $tJ\gtrsim 100$ the NESS is reached exhibiting a spatially homogeneous current with $j_{\text{ness}}=0.0225\mu J/2$.
Analogous data is presented in Fig. 3 for the case of a smaller system-bath coupling $\gamma/J=0.1$. Here, the local density in (a) as well as the local current in (b) reach the NESS at later times $tJ\gtrsim 150$ with $\nabla S_{\text{ness}}^{z}=-0.0397\mu/2$ and $j_{\text{ness}}=0.0136\mu J/2$. Note that for this case the relaxation time scale to the NESS is larger than for the previous case due to the weaker (compared to the previous case) coupling to the bath. Once again, we obtain a very reasonable agreement of data from SU and TEBD calculations.
@@
To understand the origin of the aforementioned limitation and to find a practical solution, we use the open system based method to probe the dynamical diffusion constant $\mathcal{D}_{r}(t)$, defined in Eq. (8), for a representative and fixed site $r=10$, which is not changed with $L$ and sufficiently away from the bath contact. Fixing $r$ allows for a convergence analysis with repect to time, which is presented in Fig. 5 for two different values of $\gamma/J\in\{0.1,1.0\}$ and fixed $\mu=0.1$. The insets in both panels show a zoomed-in view of the plateau region of the $\mathcal{D}_{r}(t)$ curve, highlighted by the rectangular boxes. We compute the diffusion constant for the representative site $r=10$ and $L=60$ within this plateau region at a chosen time $tJ=20$ and these results agree quite reasonably with the value of the diffusion constant extracted from the Kubo formula, as shown in the Fig. 4 (see the violet ‘$\times$’ points in panel (a)). Now, few important comments are in order: i) As can be seen in Fig. 4, these short-time results of the diffusion constant extracted from the plateau has minimal $\gamma$ dependence and as mentioned earlier, share reasonable agreement with the results from the Kubo formalism. This leads us to consider the $\gamma$ dependence of transport coefficients as a potential diagnostic of finite-size effects. ii) The diffusion constant only has a very weak time dependence within the plateau. Therefore, although changing the reference time from $tJ=20$ to some other time within the plateau would change the value of the time-dependent diffusion constant, but that change will be rather small and hence will not affect the main following conclusions for all practical purposes. On a similar note, we may point out that for $\gamma/J=0.1$, strictly speaking, $tJ=20$ is not within the ‘plateau’ for $L=20$, but the deviation from the former is also rather small and hence does not alter our main conclusions. iii) Followed by the plateau, the $\mathcal{D}_{r}(t)$ curve takes a downward turn and finally reaches the NESS. This downward turn can possibly be a signature of the onset of finite-size effects. iv) The oscillation within the plateau window emanates from the fact that TEBD is only an approximate method and additionally, the data for smaller $\gamma$ ($\gamma/J=0.1$ for example) are more ‘error prone’. This is a consequence of a weaker coupling to the baths, which is often not sufficient to suppress the operator state entanglement entropy growth (a measure of simulability) but necessary for MPS-based methods to work well [52]. Indeed we have more stable data for $\gamma/J=1.0$. v) The onset of the plateau region essentially depends on location $r$ and hence to make the comparison of different $L$ data established on a same footing, we fix the reference site. vi) As we see in the insets, the width of the plateau depends on system size $L$ and roughly scales as $\sim L$. In contrast, in a typical diffusive system, the NESS time scale behaves as $L^{2}$. This explains why the NESS results, such as $\mathcal{D}_{\text{dc}}$ obtained from open-system methods are contaminated by finite-size effects.
@@ V.2 Nonintegrable system
To this end, we employed advanced numerical techniques such as stochastic unraveling and matrix-product-state based algorithms to extract diffusion coefficients via open-system dynamics by solving the GKSL master equation. Our numerical results for dc diffusion constant $\mathcal{D}_{\text{dc}}$, extracted from the NESS, showed a mismatch with those obtained from the closed-system dynamics based on the linear-response Kubo formalism. Additionally, we observed a strong dependence of $\mathcal{D}_{\text{dc}}$ on the system-bath coupling strength $\gamma$, a phenomenon that shows up irrespective of system size considered in our work (ranging from $L=10$ to $L=60$). Such a dependence of a material constant like $\mathcal{D}_{\text{dc}}$ on details of system-bath coupling is unphysical. Therefore, our results indicate that the dc diffusion constant computed from boundary driven methods are not reliable. To find out the source of such a problem in open-system techniques and also to find a reasonable resolution, we computed the corresponding time-dependent diffusion constants $\mathcal{D}(t)$. The diffusion constant extracted from the $L$-independent plateau region of $\mathcal{D}(t)$ curve at finite times agreed well with the closed-system Kubo result and shows minimal $\gamma$-dependence. This suggests that the $\gamma$-dependence of NESS-based results may serve as a potential diagnostic for finite-size effects in the asymptotic behavior of open-system dynamics. In a similar spirit, we attributed the discrepancy in transport coefficients between open- and closed-system approaches to the incorrect order of limits inherent in open-system methods that rely on NESS quantities. Note that by definition, computing the dc diffusion constant requires taking the limits in the order $\lim_{t\to\infty}\lim_{L\to\infty}$. However, open-system numerical methods inherently reverse this order. Since these limits generally do not commute, the resulting transport coefficients usually carry finite-size errors. A rough argument is as follows: Our numerical results showed that the width of the plateau in $\mathcal{D}(t)$ scales approximately linearly with system size $L$ (for the examples studied), while the NESS time scale in a diffusive system typically scales as $\sim L^{2}$. This mismatch implies that, for any finite $L$, the NESS-based results inevitably retain finite-size effects. We emphasize that, despite limitations in computing long-time transport coefficients, the boundary-driven open system approach remains highly valuable. Beyond enabling the extraction of accurate time-dependent transport coefficients, as demonstrated in this work, it provides a powerful framework to probe transport behavior via the scaling of the steady-state current $j_{\text{ness}}$ with system size $L$. Compelling numerical evidence from previous studies [50, 49, 53, 54] showed that this method efficiently captures not only diffusive transport but also a wide range of anomalous regimes, including superdiffusion and fractonic subdiffusion.

### v2

@@ V.1 Integrable system
For this subsection, we consider the spin-1/2 XXZ chain with open boundary conditions described by Eq. (1). In Fig. 3 the real-time dynamics of (a) the local magnetization density $\langle S_{r}^{z}\rangle$ and (b) the local current $\langle j_{r}\rangle$ for a spin-$1/2$ XXZ chain with $L=20$ sites and anisotropy $\Delta=1.5$ under a strong system-bath coupling $\gamma/J=1.0$ and weak driving $\mu=0.1$ is shown. Data obtained by stochastic unraveling (SU) is averaged over $N_{\text{traj}}=50000$ trajectories. The results of both methods, SU and TEBD, show a reasonable agreement. While the remaining deviation stems from the approximative nature of the two methods (finite number of trajectories/finite bond dimension), this deviation does not affect the conclusions drawn later. As can be seen from Fig. 3, panel (a) and (b), at times $t\gtrsim 100$, the resulting NESS is reached. The NESS features a linear magnetization profile in the bulk of the system and a NESS current independent of site location $r$. As a concrete example, in Fig. 3(c), we show the corresponding magnetization profile for a chosen time $tJ=200$. By fitting over the bulk sites $r=5\dots 15$, we can extract a linear slope of $\nabla S_{\text{ness}}^{z}=-0.0421\mu/2$, which is also consistent with the TEBD calculations. In Fig. 3(b) the local current according to Eq. (7) for the bulk sites $r=5\dots 15$ is shown. As for the magnetization profile, also here the choice of the initial infinite-temperature state is reflected in $\langle j_{r}(t=0)\rangle=0$. For times $tJ\gtrsim 100$ the NESS is reached exhibiting a spatially homogeneous current with $j_{\text{ness}}=0.0225\mu J/2$.
Analogous data is presented in Fig. 4 for the case of a smaller system-bath coupling $\gamma/J=0.1$. Here, the local density in (a) as well as the local current in (b) reach the NESS at later times $tJ\gtrsim 150$ with $\nabla S_{\text{ness}}^{z}=-0.0397\mu/2$ and $j_{\text{ness}}=0.0136\mu J/2$. Note that for this case the relaxation time scale to the NESS is larger than for the previous case due to the weaker coupling to the bath. Once again, we obtain a very reasonable agreement of data from SU and TEBD calculations.
@@
To understand the origin of the aforementioned limitation and to find a practical solution, we use the open system based method to probe the dynamical diffusion constant $\mathcal{D}_{r}(t)$, defined in Eq. (8), for a representative and fixed site $r=10$, which is not changed with $L$ and sufficiently away from the bath contact. Fixing $r$ allows for a convergence analysis with respect to time, which is presented in Fig. 6 for two different values of $\gamma/J\in\{0.1,1.0\}$ and fixed $\mu=0.1$. The insets in both panels show a zoomed-in view of the plateau region of the $\mathcal{D}_{r}(t)$ curve, highlighted by the rectangular boxes. The value of the diffusion constant for the representative site $r=10$ and $L=60$ within this plateau region at a chosen time $tJ=20$ agrees quite reasonably with the value of the diffusion constant extracted from the Kubo formula, as shown in the Fig. 5 (see the violet $\times$’ points in panel (a)). Now, few important comments are in order: i) As can be seen in Fig. 5, these short-time results of the diffusion constant extracted from the plateau has minimal $\gamma$ dependence and as mentioned earlier, share reasonable agreement with the results from the Kubo formalism. This leads us to consider the $\gamma$ dependence of transport coefficients as a potential diagnostic of finite-size effects. ii) The diffusion constant only has a very weak time dependence within the plateau. Therefore, although changing the reference time from $tJ=20$ to some other time within the plateau would change the value of the time-dependent diffusion constant, but that change will be rather small and hence will not affect the main conclusions for all practical purposes. On a similar note, we may point out that for $\gamma/J=0.1$, strictly speaking, $tJ=20$ is not within the ‘plateau’ for $L=20$, but the deviation from the former is also rather small and hence does not alter our main conclusions. iii) Followed by the plateau, the $\mathcal{D}_{r}(t)$ curve takes a downward turn and finally reaches the NESS. This downward turn can possibly be a signature of the onset of finite-size effects. iv) The oscillation within the plateau window emanates from the fact that TEBD is only an approximate method and additionally, the data for smaller $\gamma$ ($\gamma/J=0.1$ for example) are more ‘error prone’. This is a consequence of a weaker coupling to the baths, which is often not sufficient to suppress the operator state entanglement entropy growth (a measure of simulability) but necessary for MPS-based methods to work well [52]. Indeed we have more stable data for $\gamma/J=1.0$. v) The onset of the plateau region essentially depends on location $r$ and hence to make the comparison between different $L$ data on a same footing, we fix the reference site. vi) As we see in the insets, the width of the plateau depends on system size $L$ and roughly scales as $\sim L$. In contrast, in a typical diffusive system, or more precisely, in the diffusive spin-$1/2$ XXZ chain, the NESS time scale behaves as $L^{2}$. This explains why the NESS results calculated from open-system approaches, such as $\mathcal{D}_{\text{dc}}$, are contaminated by finite-size effects.
@@ V.2 Nonintegrable system
To this end, we employed advanced numerical techniques such as stochastic unraveling and matrix-product-state based algorithms to extract diffusion coefficients via open-system dynamics by solving the GKSL master equation. Our numerical results for dc diffusion constant $\mathcal{D}_{\text{dc}}$, extracted from the NESS, showed a mismatch with those obtained from the closed-system dynamics based on the linear-response Kubo formalism. Additionally, we observed a strong dependence of $\mathcal{D}_{\text{dc}}$ on the system-bath coupling strength $\gamma$, a phenomenon that shows up irrespective of system size considered in our work (ranging from $L=10$ to $L=60$). Such a dependence of a material constant like $\mathcal{D}_{\text{dc}}$ on details of system-bath coupling is unphysical. Therefore, our results indicate that the dc diffusion constant computed from boundary-driven methods are not reliable.
To find out the source of such a problem in open-system techniques and also to find a reasonable resolution, we computed the corresponding time-dependent diffusion constants $\mathcal{D}(t)$, where our key findings are sketched in Fig. 1. In the long-time limit we found: (i) The dc diffusion constant $\mathcal{D}_{\text{dc}}$ extracted from the boundary-driven approach disagrees with the closed-system Kubo-based result. (ii) In the open-system setup, $\mathcal{D}_{\text{dc}}$ exhibits a strong dependence on the system-bath coupling strength $\gamma$. To identify the origin of this mismatch, we analyzed the full time evolution of $\mathcal{D}(t)$ within the boundary-driven framework and and further compare it to the Kubo-based closed-system result. Here, we found for (earlier) times $tJ\leq t_{\mathrm{fs}}J$, the open-system diffusion constant $\mathcal{D}(t)/J$ converges to a plateau value which shows a convincing agreement with the Kubo-based closed-system result and is nearly independent of system-bath coupling $\gamma$.
We suggest, that a physical interpretation follows from the finite-size time scales: For our model, the onset of finite-size effects scales as $t_{\mathrm{fs}}J\propto L$ [5]. Increasing $L$ therefore extends the initial plateau in $\mathcal{D}(t)$ and prolongs its agreement with the closed-system value, largely independent of $\gamma$. However, the approach to the NESS occurs on a (later) timescale $t^{*}J\propto L^{2}$, so NESS-based dc values from open-system approaches (such as $\mathcal{D}_{\text{dc}}$) remain affected by finite-size effects and not agreeing with the closed-system Kubo-based value. In other words, boundary-driven schemes effectively realize the order of limits $\lim_{L\to\infty}\lim_{t\to\infty}$, in contrast to $\lim_{t\to\infty}\lim_{L\to\infty}$. Since these limits generally do not commute, NESS-based results for $\mathcal{D}_{\text{dc}}$ at finite $L$ are contaminated by finite-size effects. This also allows for the interpretation of the sensitivity of $\mathcal{D}_{\text{dc}}$ to $\gamma$ from open-system approaches as a diagnostic of finite-size effects.
We emphasize that, despite limitations in computing long-time transport coefficients, the boundary-driven open system approach remains highly valuable. Beyond enabling the extraction of accurate time-dependent transport coefficients, as demonstrated in this work, it provides a powerful framework to probe transport behavior via the scaling of the steady-state current $j_{\text{ness}}$ with system size $L$. Compelling numerical evidence from previous studies [50, 49, 53, 54] showed that this method efficiently captures not only diffusive transport but also a wide range of anomalous regimes, including superdiffusion and fractonic subdiffusion.
@@ Acknowledgments
Data availability
The data that support the findings of this article are openly available at [55] .